Problem 4: Calculate the average power in a 150 resistor if the current in it is given by the waveform below.
Answer #1

I2-4 nn) time period T=3 , R=152 yoy1= nam two point formula Long = J + 11 12 13 14 dt M 10,0) 4 5 -3t ostal |-0--3-0. (t-0)imy? - 7 12 pots at X = $ [[94-at +5° 8 64-3? dt +0] Cx [+P 24-47 5 +4027, ]ox fe (la com]+, IH JE t he way =341 [1 +10+1)]